OpenStudy (anonymous):

] The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 340, how many one-step equations auditioned today? 260 340 350 670

OpenStudy (anonymous):

i need the answer as soon as possible

OpenStudy (anonymous):

plzzz

OpenStudy (anonymous):

x is the number of 1 step equations y is the number of 2 step equations x+y = 610 4x - 2y = 340 it wants you to find x

OpenStudy (anonymous):

by solving that equation

OpenStudy (anonymous):

2x+2y = 1220 + 4x-2y = 340 6x = 1560 x = 260 it is a system of equations

OpenStudy (anonymous):

the number of 1 step equations is 260

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

life saver

OpenStudy (anonymous):

you're welcome